If a train going 60 m/s hits the brakes, and it takes the train 85 seconds to stop, what is the train’s acceleration? 




1
Expert's answer
2022-03-06T18:05:43-0500

The acceleration of the train

"a=\frac{\Delta v}{\Delta t}=\frac{0-60}{85}=-0.71\:\rm m/s^2"